The number of calories in a certain dessert is normally distributed with a mean of


210 and a standard deviation of 12. Find the probability that the dessert one select will


contain between 185 and 216.

1
Expert's answer
2022-03-04T05:56:05-0500
"P(185<X<216)=P(X<216)-P(X\le185)"

"=P(Z<\dfrac{216-210}{12})-P(Z\le\dfrac{185-210}{12})"

"\approx P(Z<0.5)-P(Z\le -2.083333)"

"\approx 0.69146-0.01861\approx0.6729"